There are 37 students in an algebra class. There are 9 more girls than boys. How many girls and how many boys are in the class?
Answer by htmentor(1343)

Let B = number of boys
Then the number of girls = B + 9
The equation we need to solve is B + (B + 9) = 37
Solve for B: B = (37 - 9)/2 = 14
So the number of boys is 14 and the number of girls is 14 + 9 = 23.